### v1

Abstract
The equations for the photon surface in spherical symmetry are worked out, starting from [1], in the most general dynamical setting. We show that the condition for a timelike hypersurface to be a photon surface can be reformulated as a non-autonomous dynamical system, whose analysis reveals that the same condition also holds when the surface is generated by a null radial geodesic.
As an application, we consider a well-known model of a spherical dust cloud undergoing gravitational collapse. Comparing our findings with those in [2], we establish that the photon surface uniquely extends in the interior spacetime as a null hypersurface, allowing us to analytically investigate whether it covers the singularity developing in the LTB model.
keywords:
photon sphere, dust collapse, black hole, naked singularity
1 Introduction
Detection of black holes represents a intriguing challenge in the realm of observational astrophysics. The well-known international collaboration Event Horizon Telescope (EHT) has utilized a network of radio telescopes across the globe to elaborate an image of the black hole in galaxy M87 [3, 4], often referred to as the black hole shadow. The EHT collaboration has applied analogous methods to investigate other substantial cosmic entities also, notably the SgrA* black hole situated at the Milky Way’s core [5].
It is important to note that the EHT images are not ”direct photographs,” but rather reconstructed radio maps illustrating hot gas in orbit around the black hole, which manifest as a luminous ring.
Indeed, despite its name, what EHT aims to capture is not exactly the shape of the event horizon, but rather something that reveals the motion of objects revolving around the black hole, within the observable region to a faraway observer, which can be effectively described, in the simplest toy models, by the so-called photon sphere.
Classically, the concept of a photon sphere arises in static and spherically symmetric settings such as the Schwarzschild solution, where it manifests as a timelike hypersurface at radius $r=3M$ on which initially tangent null geodesics remain tangent as they evolve. A similar approach, based on the study of null radial geodesics, has been used in a context of static spacetimes [6, 7, 8, 9, 10, 11, 12],
When it comes to dynamical spacetimes, the situation is more involved. One could think to generalize the above approach to this situation, which actually proves fruitful where the metric assumes a special form - for example, [13, 14, 15] study the situation for a spherical metric written in Bondi coordinates $(v,r,\theta,\phi)$, where $v$ is a null coordinate, as is the case for the Vaidya solution.
However, this notion has been extended and generalized to a broader geometric object known as the photon surface, formalized in the seminal work by Claudel, Virbhadra, and Ellis [1], where it is defined independently of any specific symmetry or Killing structure.
In their framework, photon surfaces are understood as nowhere-spacelike hypersurfaces invariant under the flow of null geodesics that remain confined within the surface. This perspective allows for a more flexible analysis, applicable not only in static spacetimes [16] but also in dynamical regimes where symmetry assumptions are weakened or absent. Subsequent studies, such as those by Cao and Song [2], have proposed quasi-local characterizations of photon surfaces in general spherically symmetric spacetimes, aiming to preserve the spirit of Claudel et al.’s geometric definition while adapting it to the complexity of evolving gravitational systems.
One physically rich setting in which photon surfaces can be investigated dynamically is gravitational collapse. The Lemaitre–Tolman–Bondi (LTB) models, describing inhomogeneous spherical dust clouds undergoing collapse under their own gravity, provide a valuable laboratory for probing how such surfaces behave in the presence of singularities and horizons. These models allow for analytical control over the metric and matter profiles and, depending on the initial data, can yield either black hole formation or naked singularities, offering insight into the end-state of light rays and causal curves near regions of extreme curvature [17]. Recent investigations have demonstrated that shadows and photon spheres, traditionally associated with black holes, can also arise in certain naked singularity spacetimes resulting from gravitational collapse, depending on specific parameters of the model [18]. This challenges the conventional interpretation of shadow observations as definitive evidence for black holes.
The central aim of this work is to explore how photon surfaces, originally defined in static contexts, can be meaningfully extended into dynamical spacetimes—specifically, within the LTB collapse model. By deriving and analyzing the governing equations for photon surfaces in the most general spherically symmetric dynamical setting, we assess whether and how these surfaces can be propagated from the exterior Schwarzschild geometry into the interior region occupied by the collapsing matter. In particular, we focus on whether such an extension can effectively cover the singularity forming at the center, depending on the visibility of that singularity.
To carry out this investigation, we start by reviewing the photon surface condition for timelike hypersurfaces in spherically symmetric spacetimes, as given in [1], and derive the relevant differential equations that such surfaces must satisfy. We then apply this framework to the LTB model, examining how the photon surface extends across the boundary between interior and exterior spacetimes and determining the precise conditions under which such an extension exists and remains physically meaningful. In this respect, the present paper aims to compare and discuss its findings with those presented in [2], trying to highlight some aspects that could be otherwise underestimated, particularly regarding the relation between the photon surface behavior and the appearance of naked singularities.
The paper is organized as follows. In Section 2, we recall the geometric setting for photon surfaces in spherically symmetric spacetimes and derive their evolution equations. Section 3 introduces the marginally bound LTB collapse model and discusses how the photon surface condition adapts to this context. We analyze the possible extensions of the photon surface from the Schwarzschild exterior into the collapsing interior, distinguishing between spacelike extensions and those corresponding to null radial geodesics. In Section 4, we focus on the qualitative behavior of these surfaces, examining their ability to shield the central singularity. We show that the nature of the singularity—naked or covered—dictates whether the photon surface terminates at the regular center or reaches the singularity itself. Finally, Section 5 summarizes the main findings and suggests possible directions for further investigation.
2 Photon surfaces in spherical symmetry
It is a well-established fact that Schwarzschild spacetime exhibits $r=3m$ as a hypersurface with the property that initially tangent null geodesics remain tangent, called photon sphere. As we shall see, the paper [1] suggests that photon spheres arise as particular cases of more general objects, photon surfaces.
Definition 1.
A photon surface of $(M,g)$ is an immersed, nowhere–spacelike hypersurface $S$ of $(M,g)$ such that $\forall p\in S$ and $\forall k\in T_{p}S$, $\exists\gamma:(-\epsilon,\epsilon)\to M$ such that $\gamma(0)=p,\,\dot{\gamma}(0)=k$, and $\gamma((-\epsilon,\epsilon))\subset S$.
Remark 1.
A null hypersurface is trivially a photon surface.
In the spherically symmetric case, [1] states a condition for a timelike hypersurface to be a photon surface:
Theorem 1.
[1, Theorem III.1] Let $(M,g)$ be a spherically symmetric spacetime and $S$ an SO(3)SO(3)–invariant timelike hypersurface of $(M,g)$ and let $X$ be the SO(3)SO(3)–invariant unit future–directed timelike tangent vector field along $S$ orthogonal to the SO(3)SO(3)–invariant 2–spheres in $S$. Let $\mathcal{T}$ be one such SO(3)SO(3)–invariant 2–sphere in $S$ and $\mathcal{T}_{s}$ be the SO(3)SO(3)–invariant 2–sphere in $S$ at arc length $s$ from $\mathcal{T}$ along the integral curves of $X$. Then $S$ is a photon surface of $(M,g)$ iff
$\frac{\mathrm{d}^{2}}{\mathrm{d}s^{2}}A_{s}=\frac{1}{4A_{s}}\left(\frac{\mathrm{d}}{\mathrm{d}s}A_{s}\right)^{2}+A_{s}\left(\frac{1}{3}\Theta^{2}-G_{\alpha\beta}n^{\alpha}n^{\beta}\right)-4\pi,$ (1)
where $A_{s}$ is the area of $\mathcal{T}_{s}$ and $\Theta$ is the expansion of the normal unit $n^{\alpha}$ to $S$.
Now, let us consider a spacetime $M$ with spherical symmetry, characterized by coordinate variables $(t,x,\theta,\phi)$. The metric is expressed as follows:
$g=-e^{2\nu}\,\mathrm{d}t^{2}+e^{2\lambda}\,\mathrm{d}x^{2}+r^{2}\mathrm{d}\Omega^{2},$ (2)
where $\lambda$, $\nu$, and $r$ are functions of $(t,x)$ and $\mathrm{d}\Omega^{2}=\mathrm{d}\theta^{2}+\sin^{2}\theta\,\mathrm{d}\phi^{2}$. We shall proceed to apply Theorem 1. Since $S$ is timelike, we can parameterize it with coordinates $(t,x(t),\theta,\phi)$, with $x(t)$ to be determined from the equation (1). In this situation, indeed, the area of $\mathcal{T}_{s}$ is given by $A_{s}=4\pi x(t(s))$, and (1) takes the form
$\ddot{x}(t)=\frac{\left(r_{x}-r\nu_{x}\right)e^{2(\nu-\lambda)}}{r}+\dot{x}(t)\left(\frac{r_{t}}{r}-2\lambda_{t}+\nu_{t}\right)\\ +\dot{x}(t)^{2}\left(-\frac{r_{x}}{r}-\lambda_{x}+2\nu_{x}\right)+\dot{x}(t)^{3}\frac{\left(r\lambda_{t}-r_{t}\right)e^{2(\lambda-\nu)}}{r},$ (3)
where $\dot{x}(t)$ and $\ddot{x}(t)$ stand for derivatives of $x$ with respect to $t$ and, given a function $f(t,x)$, $f_{t}$ and $f_{x}$ denote its partial derivatives. We can transform this second-order ODE into a first-order system of ODEs as follows
$\dot{x}(t)$ $=e^{\nu-\lambda}y(t),$ (4a)
$\dot{y}(t)$ $=\left(1-y(t)^{2}\right)\frac{e^{\nu-\lambda}\left(r_{x}-r\nu_{x}\right)+y(t)\left(r_{t}-r\lambda_{t}\right)}{r}.$ (4b)
Remark 2.
Inspection of (4a)–(4b) above leads us to conclude that, if $y(t_{0})=\pm 1$ for some $t_{0}$ then $y(t)=\pm 1$ throughout the solution. Let us notice that $\dot{x}=\pm e^{\nu-\lambda}$ is the equation of a radial null geodesic in this spacetime. Therefore, although we already know (see Remark 1) that null hypersurfaces are trivially photon surfaces, we can see that equation (1) in Theorem 1 – that in principle assumes the surface to be timelike – also holds for those null hypersurfaces $S$ generated by a radial null geodesic $(t,x(t))$ in the 2–dimensional quotient submanifold $M/S^{2}$. We should remember this fact, which will prove helpful when discussing the initial conditions for equation (3) or, equivalently, the system (4a)–(4b). As we will observe in the examples we explore later, these initial conditions are essentially determined by the boundary conditions applied to the metric being analyzed.
Example 1.
Let us consider the particular case where (2) is static, therefore admitting a $SO(3)\times\mathbb{R}$ group of isometries, where the $\mathbb{R}$ isometries are generated by a Killing vector field $K$, orthogonal to the $SO(3)$ orbits. In this context, the notion of photon sphere is defined in [1] as an $SO(3)\times\mathbb{R}$ invariant photon surface. In the static case, $\nu,\lambda$ and $r$ are functions of $x$ only, and the Killing vector field that generates the isometries of $\mathbb{R}$ is collinear to $\partial_{t}$. Therefore, a photon space is a photon surface with constant $x(t)$, which in view of (4a)–(4b) is implicitly defined by those $x$ satisfying
$\nu^{\prime}(x)=\frac{r^{\prime}(x)}{r(x)}.$ (5)
For instance, in the case of the anisotropic generalizations of de Sitter spacetimes [19] given by  [^1] Please note that in this example we are using $r$ as a coordinate in place of $x$, in compliance with most of the literature on the subject.
$g=-\left(1-\frac{2m(r)}{r}\right)\,\mathrm{d}t^{2}+\left(1-\frac{2m(r)}{r}\right)^{-1}\,\mathrm{d}r^{2}+r^{2}\mathrm{d}\Omega^{2},$ (6)
equation (5) takes the well-known form (see [1, (67)])
$1-\frac{3m(r)}{r}=m^{\prime}(r),$
which in the special Schwarzschild case $m$ constant is solved exactly by $r=3m$.
3 Photon surfaces in dust collapse
Now, let us apply the notions sketched in the above section to the very well-known model of a spherical dust cloud undergoing gravitational collapse. This is modeled by considering a general spherical metric (2) where the energy–momentum tensor is given by $T=-\rho\,\mathrm{d}t\otimes\partial_{t}$ [20, 17]. The metric is given by the Lemaître–Tolman–Bondi (LTB) model
$g=-\mathrm{d}t^{2}+\frac{r_{x}^{2}}{1+k(x)}\,\mathrm{d}x^{2}+r^{2}\,\mathrm{d}\Omega^{2},$ (7)
and the collapse evolution is dictated by the PDE
$r_{t}(t,x)=-\sqrt{k(x)+\frac{2m(x)}{r(t,x)}}.$ (8)
The initial data $k(x)$ and $m(x)$ are prescribed at some initial time $t_{0}$, i.e., $t_{0}=0$, together with the initial energy $\rho_{0}(x):=\rho(t_{0},x)$. The value of the Misner-Sharp mass $m(x)=\frac{r}{2}\left(1-g^{\alpha\beta}r_{\alpha}r_{\beta}\right)$ will also determine the initial condition $r(t_{0},x)$, because integrating the Einstein field equation $m_{x}=4\pi\rho r^{2}r_{x}$ along $t=t_{0}$ one obtains
$r(t_{0},x)^{3}=\int_{0}^{x}\frac{3m^{\prime}(\xi)}{4\pi\rho_{0}(\xi)}\,\mathrm{d}\xi.$
As pointed out in [2, (5.5)], for the LTB model a lengthy but straightforward computation transforms the equation (3) for photon surfaces as
$\ddot{x}(t)=\frac{k(x)+1}{rr_{x}}+\left(\frac{r_{t}}{r}-\frac{2r_{tx}}{r_{x}}\right)\dot{x}(t)\\ -\left(\frac{r_{x}}{r}+\frac{r_{xx}}{r_{x}}-\frac{k^{\prime}(x)}{2(k(x)+1)}\right)\dot{x}(t)^{2}+\frac{r_{x}\left(r\,r_{tx}-r_{x}r_{t}\right)}{r(k(x)+1)}\dot{x}(t)^{3},$ (9)
which can be given in the form of the 1st-order system (4a)–(4b):
$\dot{x}(t)$ $=\frac{\sqrt{k(x)+1}}{r_{x}}\,y(t),$ (10a)
$\dot{y}(t)$ $=(1-y(t)^{2})\frac{r_{x}\sqrt{k(x)+1}+y(t)\left(r_{x}r_{t}-r\,r_{tx}\right)}{r\,r_{x}}.$ (10b)
When it comes to assigning initial conditions for this system, we have first to observe that the metric (7) describes only the interior part of the model we are describing, i.e. the dust star. In other words, this metric is defined for $x\in[0,x_{b}]$, where $x_{b}$ represents the boundary of the star. The timelike hypersurface $\Sigma=\{x=x_{b}\}$ must be matched with a suitable exterior solution that describes the matter distribution outside the spherical dust cloud. It is well known that for the exterior region to be Schwarzschild spacetime, the radial pressure, indicated by the $T^{1}_{1}$ component of the energy-momentum tensor, must be zero at $\Sigma$. According to the Einstein field equations, this condition ensures that the Misner–Sharp mass does not depend on time $t$ and remains constant across the matching hypersurface, a necessary condition for matching [21]. Naturally, this applies to the model being analyzed here, where pressures uniformly vanish throughout the spacetime.
From now on we will confine ourselves to the so-called marginally bound case, which amounts to set $k(x)=0$. This choice does not exclude any of the physical implications of the model, as one can demonstrate [22]; however, it allows the equation (8) to explicitly integrate and yield
$r(t,x)=r(1-\mu(x)t)^{2/3},\qquad\mu(x)=\frac{3}{2}\sqrt{\frac{2m(x)}{x^{3}}},$ (11)
where also the initial profile of energy density has been fixed in such a way that $r(0,x)=x$. In this way, each timelike hypersurface with fixed $x=x_{0}$ physically represents the shell of dust cloud particles that, at initial time, are located at a proper distance $x_{0}$ from the center of symmetry, and the function $t\mapsto r(t,x_{0})$ will represent the dynamical evolution of this shell as time flows by. The shell collapses under its own gravitational influence and eventually becomes singular at comoving time $t_{s}(x_{0})=\mu(x_{0})^{-1}$, where the energy density $\rho$ diverges, indicating the presence of a spacetime singularity, which can be covered by the presence of the apparent horizon $t_{h}(x_{0})$, described by the equation $r(t_{h}(x),x)=2m(x)$, see Figure 1.
[…]
Figure 1: Schematic picture of the collapse of spherical dust. The process begins from a regular configuration at $t=0$. Every shell $x\in[0,x_{b}]$ within the spherical cloud undergoes collapse, becoming trapped at comoving time $t_{h}(x)$ and fully collapsing at $t_{s}(x)$. The symmetry center gets trapped simultaneously as it becomes singular, allowing for the chance that photons might escape from the central singularity and reach the region $r>2m$.
We will refer to this model as the marginally bound LTB collapse. Its endstate, starting from regular initial conditions at $t=0$, is well known in the literature [20]: let us summarize it briefly. Consider a mass profile such that, using (11)
$\mu(x)=1-ax^{n}+o(x^{n}),$ (12)
with $a>0$ and $n\in\mathbb{N}\setminus\{0\}$: in this way the initial energy $\rho_{0}(x)$ is regular up to the central shell $x=0$, and decreasing in a right neighborhood of $x=0$ – we can suppose that the star boundary $x_{b}$ belongs to this neighborhood. In this configuration, the non-central singularities, i.e. $t_{s}(x)$ for every $x\in]0,x_{b}]$ are always covered by the apparent horizon. Conversely, the central singularity $t_{s}(0)$ may be either covered or naked, depending on the choice of parameters $a$ and $n$ in (12). Indeed, one has to observe that deriving $t_{h}(x)$ from the condition $r=2m$ leads to the expression
$t_{h}(x)=\frac{1}{\mu(x)}\left(1-\frac{8}{27}\mu(x)^{3}x^{3}\right),$
implying $t_{h}(0)=t_{s}(0)=1$, see again Figure 1.
Full details of these studies are based on the qualitative study of radial null geodesics, and can be found in [23, 22]. According to these findings, the central singularity is:
  • naked if $n=1,2$ or $n=3$ and $a>\frac{2(26+15\sqrt{3})}{27}$, and
  • covered, otherwise.
To sum up, the global model that we are going to consider will therefore be composed of the interior LTB metric (7) with $k(x)=0$ and $r(t,x)$ given by (11), defined in the set
$\{(t,x)\,:\,x\in[0,x_{b}],\,\,t<t_{s}(x)\},$
matched to a Schwarzschild exterior
$g_{E}=-\left(1-\frac{2M}{r}\right)\,d\tau^{2}+\frac{1}{1-\frac{2M}{r}}\,\mathrm{d}r^{2}+r^{2}\,\mathrm{d}\Omega^{2}$ (13)
at the star boundary $\Sigma=\{x=x_{b}\}$. Israel–Darmois junction conditions [24] guarantee that the matching is $C^{1}$ across $\Sigma$, provided that the Schwarzschild mass is given by $M=m(x_{b})$. Observe that the coordinates $(\tau,r)$ in (13) will be related to $t$ on the matching surface – in particular, $r=r(t,x_{b})$ on $\Sigma$ [^2] That justifies the slight abuse of notation where we employ the same letter $r$ for both the radial coordinate in (13) and the metric coefficient (11) within (7).. Details can be found in [25].
Let us now consider the photon surface for this model. On the exterior (13), of course, the photon surface is given by $r=3M$. When it comes to determining its extension to the interior spacetime, we need to set initial conditions for the system (10a)–(10b). The first condition is necessarily given by $x(t_{0})=x_{b}$, where $t_{0}$ satisfies the equation $r(t_{0},x_{b})=3m(x_{b})$, i.e., using (11),
$t_{0}=\frac{1}{\mu(x_{b})}\left(1-(2/3)^{3/2}\mu(x_{b})^{3}x_{b}^{3}\right)=\frac{\sqrt{2}}{3}\left(\sqrt{\frac{x_{b}^{3}}{m(x_{b})}}-3\sqrt{3}m(x_{b})\right)$ (14)
Consequently, the second condition must be expressed as $y(t_{0})=y_{0}$. In the following, we are going to show that $y_{0}=1$ is the only physically acceptable option. Let’s see why.
First of all, in order for the extension not to be spacelike at $t_{0}$, we need $|y_{0}|\leq 1$, as we can see using (10a) and (7).
In view of Remark 1 we can accept $y_{0}=\pm 1$ as initial conditions, since they will correspond to the outgoing ($y_{0}=1$) and ingoing ($y_{0}=-1$) radial null geodesics starting at $x=x_{b}$. Among the above cases, the outgoing curve will necessarily cover the ingoing one, which makes the $y_{0}=1$ extension more significant than the other.
Of course, one may consider other extensions which start spacelike at $x=x_{b}$. These choices would correspond to setting $|y_{0}|<1$. However, $y_{0}<0$ would correspond to a decreasing $x(t)$ curve, and for the same reason would be less significant than the extension that we have considered with $y_{0}=1$. The following lemma will help us rule out all the other cases.
Lemma 1.
Let $(x(t),y(t))$ solution of (10a)–(10b) such that $(x(t_{0}),y(t_{0}))=(x_{b},y_{0})$ with $y_{0}\in]0,1[.$ Then $\exists\tilde{t}<t_{0}\,:\,x(\tilde{t})=x_{b}.$
Proof.
Before presenting the proof, Figure 2 illustrates the core concept it is based on: the (non–autonomous) vector field induced by (10a)–(10b) prevents the existence of curves with initial data $|y_{0}|<1$ that extend back to $x=0$.
Figure 2: The behavior of the vector field associated to (10a)–(10b) is such that the only way to left–extend a solution from $x=x_{b}$ to $x=0$ is when $y_{0}=1$. Notice that the field is non autonomous (here is represented the situation at $t=1.01$ only). Other parameters are n=2,a=1.5n=2,\,a=1.5.
After the above intuitive idea, let us start the proof by showing, as an intermediate step, that $\exists t_{1}<t_{0}\,:\,y(t_{1})=0$. By contradiction, suppose that $y(t)>0,\forall t<0$. Then $x(t)$ is increasing $\forall t<t_{0}$ and then, called $t_{\text{inf}}\in\mathbb{R}\cup\{+\infty\}$ the infimum of the maximal interval where the solution can be extended, there exists $\lim_{t\to t_{\text{inf}}}x(t)=x_{0}\in[0,x_{b}[$.
Suppose $t_{\text{inf}}=$-$\infty$. Using (11) and (12) in (10a)–(10b) we get
$\frac{\dot{y}(t)}{1-y(t)^{2}}\sim\frac{1}{x(1-ax^{n})^{2/3}}\cdot\frac{1}{t^{2/3}},\quad\text{\ as\ }t\to-\infty,$ (15)
which implies the existence of positive constants $C_{1}$, $C_{2}$ such that
$\frac{1-y(t)}{1+y(t)}\geq e^{-C_{1}t^{1/3}+C_{2}}.$
Since the right-hand side positively diverges as $t\to$-$\infty$, we get a contradiction, recalling that $y(t)$ should remain positive. Then $t_{\text{inf}}\in\mathbb{R}$, and this situation takes place only if $\lim_{t\to t_{\text{inf}^{+}}}x(t)=0$. But in this case, again (10a)–(10b) gives
$\frac{\mathrm{d}y}{\mathrm{d}x}=\frac{\dot{y}}{\dot{x}}\sim\frac{1-y^{2}}{x\,y},\quad\text{\ as\ }x\to 0$ (16)
from which one gets the existence of a positive constant $C_{3}$ such that
$1-y(x)^{2}\sim\frac{C_{3}}{x^{2}}.$
Again, the right-hand side above positively diverges, which is excluded here.
Then so far we have shown that $y(t_{1})=0$ for some $t_{1}<t_{0}$. Inspection of the equivalent equation (9) shows that if $\dot{x}=0$, then $\ddot{x}>0$, then $x(t_{1})$ is a local minimum for $x(t)$, and so in a left neighborhood of $t_{1}$, $x(t)$ is decreasing and then $y(t)$ is negative. This fact forces $y(t)<0,\,\forall t<t_{1}$ and then $x(t)$ will be decreasing for earlier times than $t_{1}$. Let us show that $x(\bar{t})=x_{b}$ for some $\bar{t}<t_{1}$. If that is not the case, then $(x(t),y(t))$ will remain in an area where the ODE is regular ($x(t)>x(t_{1})$ and $y\in]-1,0]$). This means that the solution can be extended $\forall t<t_{1}$. But in this case the same estimate as (15) gives $y(t)\to-1$, and using (10a) we get the existence of a positive constant $C_{4}$ such that
$\dot{x}\sim-\frac{C_{4}}{t^{2/3}},$ (17)
that integrated as $t\to$-$\infty$ gives $x(t)$ positively diverging, which is a contradiction, and this concludes the proof of the lemma. ∎
The above Lemma shows some extension of photon surfaces that are not physically acceptable. The reason is that these extensions have two points of contact with the star boundary - one at $t_{0}$ and the other one at an earlier comoving time $\bar{t}<t_{0}$. Since we are considering a collapsing model, i.e. $r_{t}<0$, we have $r(\bar{t},x_{b})>r(t_{0},x_{b})=3M$. Therefore, the extension at $\bar{t}$ cannot be matched with the exterior photon space, that indeed is matched at $t=t_{0}$. We conclude that the null extension is the only possible one: the findings of this section can be then summarized in the following.
Theorem 2.
The marginally bound LTB collapse admits an extension of the photon surface $r=3M$ in the interior of the cloud. This is the only physically acceptable extension, and it is given by the surface
$\{(t,x(t),\theta,\phi)\,:\,t<t_{0}\},$
with the property that, fixed $\theta$ and $\phi$, the corresponding curve is a null radial geodesic, i.e. the couple $(x(t),1)$ solves system (10a)–(10b) with $k(x)\equiv 0$, given initial data $x(t_{0})=x_{b}$ and $y(t_{0})=1$.
Remark 3.
The initial condition $x(t_{0})=x_{b}$ ensures that $m$ is $C^{0}$ along the boundary of the star. Naturally, the Misner-Sharp mass remains constant on the exterior solution, and consequently, on the exterior photon surface at $r=3M$. Therefore, demanding that $m$ possesses $C^{1}$ regularity at the star’s boundary effectively means requiring
$\frac{\mathrm{d}}{\mathrm{d}t}m(x(t_{0}))=0,$
which is claimed in [2] to be assumed  [^3] Observe that in [2] Misner-Sharp mass is denoted by $E$. as a second initial condition, besides $x(t_{0})=x_{b}$. But actually, that would give $\dot{x}(t_{0})=0$ since the mass is a function of $x$ only and in fact, from the second order ODE (9) we have that, when $\dot{x}(t)=0$ then $\ddot{x}(t)=1/(r\,r_{x})>0$ and so a critical point of the curve $x(t)$ is necessarily a strict local minimum. That excludes this possibility because, if we search for an extension of the photon surface in the interior portion of the model, we would rather look for a function $x(t)<x_{b}$, and this cannot happen in this case where, in a neighborhood of the initial time $t_{0}$, $x(t)$ would attain strictly greater values than $x(t_{0})=x_{b}$.
The apparent contradiction is indeed resolved because, as a matter of fact, the condition used in [2] is precisely that coming from our Theorem 2, as one can verify looking at [2, (5.41)] – although misinterpreted there, as explained above.
4 Photon surface analysis
Once we have established that the photon surface uniquely extends in the interior spacetime as a null hypersurface, let us study its qualitative behavior, to understand whether it covers the singularity developing in the LTB model.
For this analysis, it will be more useful to denote the photon surface as a solution of the outgoing null radial geodesic expressing $t$ as a function of $x$, i.e.
$\frac{\mathrm{d}t(x)}{\mathrm{d}x}={r_{x}(t(x),x)},\qquad t(x_{b})=t_{0},$ (18)
and defined in a left neighborhood of $x_{b}$, possibly extendable up to the regular centre $x=0$. That is possible because $r_{x}>0$ and therefore $x(t)$ can be inverted. In this way we can exploit all the techniques known in literature and based on comparison theorems in ODE [23, 22]. These techniques are based on a remarkable property [26] of the apparent horizon curve $t_{h}(s)$ to be a subsolution of the null geodesic equation $t^{\prime}(x)=r_{x}(t(x),x)$, i.e.
$t_{h}(x)\leq r_{x}(t_{h}(x),x).$
Recalling that $t_{h}(x_{b})>t_{0}$, we have that $t(x)<t_{h}(x)$, $\forall x<x_{b}$. Please observe that there still might be the case that $t(0)=t_{h}(0)=1$, because on the singular center of symmetry $(r=0,t=1)$ the regularity needed to apply comparison theorems in ODE is lost.
Then, the photon surface will cover the singularity completely (i.e. including the center) if one of the two following instances happens:
  1. 1.
    $\lim_{x\to 0^{+}}t(x)<t_{s}(0)=1$, or
  2. 2.
    $\lim_{x\to 0^{+}}t(x)=t_{s}(0)=1$ but there are no null radial outgoing geodesics $t_{g}(x)$ such that $t_{g}(0)=1$ and $t_{g}(x)<t(x)$, $\forall x\in]0,x_{b}]$.
We will observe that the situation varies depending on whether the central singularity is naked or not. The initial scenario arises when, in (12), we have $n=1,2$ or $n=3$ and $a>{2(26+15\sqrt{3})}/{27}$. These cases specifically align with those where we can identify a super-solution for (18) of the form $t_{\xi}=1+\xi x^{n}$, selecting $\xi$ such that $t_{\xi}(x)<t_{h}(x)$ in a right neighborhood of $x=0$. The star boundary $x_{b}$ will be chosen in this neighborhood.
Assume that it is possible to select $t_{0}$ as defined in (14) such that $t_{\xi}(x_{b})<t_{0}<t_{h}(x_{b})$. In this case, the continuation of the solution to the Cauchy problem (18) extends to smaller values of $x$, reaching $x=0$, which results in $t(x)\to 1$. Consequently, the extension of the photon surface originates from the central singularity. Moreover, consider a time $t^{*}\in]t_{\xi}(x_{b}),t_{0}[$, the radial null geodesic with initial condition $t(x_{b})=t^{*}$ will be extended backwards up to $x=0$, again with $t(x)\to 1$, giving rise to a photon (actually, an infinite number of photons, given the arbitrariness of $t^{*}$) emanating from the central singularity but escaping the photon surface.
In view of the above, and recalling that
$r(t_{h}(x_{b}),x_{b})=2m(x_{b})<3m(x_{b})=(t_{0},x_{b}),$
we only have to check that $t_{\xi}(x_{b})<t_{0}$. Using (14) we obtain that this condition is satisfied in a right neighborhood of $x=0$ if and only if $n=1,2$ or $n=3$ and $a>(2/3)^{3/2}$. Therefore the mass profiles ensuring a central naked singularity also imply that the photon surface extends up to the central singularity, but it is insufficient to cover it.
The situation described here is represented in Figures 3 and 4, where the parameters are set in order to have a naked central singularity.
Figure 3: The extension of the photon surface in case when the central singularity is naked. The parameters of (12) are set as follows: $n=2$, $a=1.5$, with the star boundary $x_{b}=0.33$. The curve $t_{\xi}(x)$, with $\xi=1.35$, is a supersolution of the ODE (18) in $0,x_{b}]$. The photon surface is extended back to the central singularity.
Figure 4: Another extension of the photon sphere with a central naked singularity. Here $n=3$, $a=8$ and $x_{b}=0.3$. The parameter of the supersolution here is $\xi=6$.
The curve $t_{\xi}(x)$ is a supersolution of (18) and $x_{b}$ chosen such that $t_{0}\in]t_{\xi}(x_{b}),t_{h}(x_{b})[$. The solution of (18) can be extended back for $x<x_{b}$: it cannot cross neither the subsolution $t_{h}(x)$ from below nor the supersolution $t_{\xi}(x)$ from above, then it must extend up to the central singularity $(x=0,t=1)$.
Conversely, when $n>3$ or $n=3$ with a sufficiently small $a$, resulting in a covered central singularity, the photon surface will extend back to the regular center, see Figure 5.
Figure 5: The extension of the photon sphere, in case the central singularity is not naked, is necessarily continued back to the regular centre $t<1$. Here $n=4$, $a=8$, $x_{b}=0.3$.
If it were to reach the central singularity, it would be a null radial geodesic escaping the central singularity, violating the conditions as the central singularity itself is trapped in this context.
We can summarize the above discussion in the following
Theorem 3.
In the marginally bound LTB collapse, the photon surface extends back to the central singularity if and only if this is naked. In this case, there exist null radial geodesics $t_{g}(x)$ emanating from the central naked singularity such that $t_{g}(x)<t(x)$, $\forall x\in]0,x_{b}]$, where $t(x)$ is the solution of (18) associated with the photon surface extension established in Theorem 2.
Otherwise, if the central singularity is covered, the photon surface will extend to the regular centre.
5 Discussion and conclusions
The study of photon surfaces has generated increasing attention in recent years, particularly due to their fundamental role in the causal structure and observational features of spacetimes. In this work, we investigated the behavior and possible extensions of photon surfaces within a dynamically evolving spherically symmetric spacetime, particularly in the context of the marginally bound Lemaitre-Tolman-Bondi (LTB) dust collapse model. Building on the general framework for photon surfaces established in earlier literature, we derived the dynamical equations governing their evolution and showed that, under suitable initial conditions, these surfaces can be uniquely continued into the interior of the collapsing matter as null hypersurfaces. This analysis not only confirms but also refines previous claims, clarifying that the condition used in earlier works—interpreted there as a requirement for the continuity of the mass profile—is more accurately understood in terms of the geometric nature of null hypersurfaces.
Our results indicate that the only physically consistent extension of the photon surface $r=3M$ into the collapsing cloud is given by a surface generated by outgoing null radial geodesics. This extension satisfies the photon surface equation and adheres to the junction conditions at the boundary with the Schwarzschild exterior. Importantly, the nature of the central singularity—whether it is covered or naked—plays a crucial role in determining the global structure of the photon surface. When the singularity is covered, the photon surface terminates at the regular center before the singularity forms. Conversely, if the singularity is naked, the photon surface can extend back all the way to the central singularity itself. However, in such cases, the surface is not sufficient to prevent the escape of photons from the singularity, highlighting its limited covering capability.
These findings, relying on purely analytical studies, refine previous contributions – in particular [2] – and contribute to the ongoing discussion about the role of photon surfaces in dynamical spacetimes and their potential connection to cosmic censorship, showing that photon surfaces, though geometric in nature, are deeply intertwined with the causal and global structure of the underlying spacetime. Future research may explore whether similar conclusions hold in more general collapse scenarios, including non-marginally bound cases or models with non-zero pressure, as well as in the presence of rotation or other deviations from spherical symmetry.